Prove that following vectors are perpendicular to each other

(i) ๐‘จโƒ—โƒ— =(๐’Šฬ‚+๐Ÿ๐’‹ฬ‚+๐Ÿ‘๐’Œฬ‚) and ๐‘ฉโƒ—โƒ— =(๐Ÿ๐’Šฬ‚โˆ’๐’‹ฬ‚)

(ii) ๐‘ชโƒ—โƒ— =(๐Ÿ‘๐’Šฬ‚โˆ’๐Ÿ๐’‹ฬ‚+๐Ÿ–๐’Œฬ‚) and ๐‘ซโƒ—โƒ— =(๐Ÿ๐’Šฬ‚+๐Ÿ๐’‹ฬ‚โˆ’๐Ÿ/๐Ÿ’๐’Œฬ‚)

(iii) ๐‘ฌโƒ—โƒ— =(๐Ÿ๐Ÿ–๐’Šฬ‚+๐Ÿ”๐’‹ฬ‚+๐Ÿ”๐’Œฬ‚) and ๐‘ญโƒ—โƒ— =(โˆ’๐Ÿ๐’Šฬ‚+๐Ÿ‘๐’‹ฬ‚+๐Ÿ‘๐’Œฬ‚)

(iv) ๐‘ฎโƒ—โƒ— =(โˆ’๐Ÿ”๐’Šฬ‚+๐Ÿ‘/๐Ÿ’๐’Œฬ‚) and ๐‘ฏโƒ—โƒ—โƒ— =(๐Ÿ/๐Ÿ๐’Šฬ‚+๐Ÿ/๐Ÿ‘๐’‹ฬ‚+๐Ÿ’๐’Œฬ‚)


1
Expert's answer
2021-03-09T10:50:18-0500

i) False


"\begin{vmatrix}
 i & j & k\\
 1 & 2 & 3\\
 2 & -1 & 0
\end{vmatrix}=(3,6,-5)"

ii) False


"\begin{vmatrix}
 i & j & k\\
 3 & -2 & 8\\
 2 & 2 & -0.25
\end{vmatrix}=(-15.5,16.75,10)"

iii) False


"\begin{vmatrix}
 i & j & k\\
 18 &6 &6\\
 -2 & 3 & 3
\end{vmatrix}=(0,-66,66)"

iv) False


"\begin{vmatrix}
 i & j & k\\
 -6 &0 &0.75\\
 0.5 & \frac{2}{3} & 4
\end{vmatrix}=(-0.5,24.375,-4)"